# FAST-GSC: Fast and Adaptive Semantic Transmission for Generative Semantic Communication

### v1

@@ Abstract
The rapidly evolving field of generative artificial intelligence technology has introduced innovative approaches for developing semantic communication (SemCom) frameworks, leading to the emergence of a new paradigm—generative SemCom (GSC). However, the complex processes involved in semantic extraction and generative inference may result in considerable latency in resource-constrained scenarios. To tackle these issues, we introduce a new GSC framework that involves fast and adaptive semantic transmission (FAST-GSC). This framework incorporates one innovative communication mechanism and two enhancement strategies at the transmitter and receiver, respectively. Aiming to reduce task latency, our communication mechanism enables fast semantic transmission by parallelizing the processes of semantic extraction at the transmitter and inference at the receiver. Preliminary evaluations indicate that while this mechanism effectively reduces task latency, it could potentially compromise task performance. To address this issue, we propose two additional methods for enhancement. First, at the transmitter, we employ reinforcement learning to discern the intrinsic temporal dependencies among the semantic units and design their extraction and transmission sequence accordingly. Second, at the receiver, we design a semantic difference calculation module and propose a sequential conditional denoising approach to alleviate the stringent immediacy requirement for the reception of semantic features. Extensive experiments demonstrate that our proposed architecture achieves a performance score comparable to the conventional GSC architecture while realizing a 52% reduction in residual task latency that extends beyond the fixed inference duration.
Index Terms:
Semantic Communication, Generative AI, Prompt Engineering, Diffusion
@@ I Introduction
The imminent advent of sixth-generation (6G) networks is anticipated to experience an unparalleled increase in data traffic, fueled by the proliferation of cutting-edge applications such as ultra-high-definition video streaming, extended reality, and extensive Internet of Things (IoT) deployments. Concurrently, traditional communication systems are approaching their limit, and the scarcity of available spectrum resources is becoming more pronounced. These factors collectively present substantial challenges in ensuring the quality of user experiences for these advanced services [1]. Recognizing the capacity of generative artificial intelligence (GAI) for modal transformation and its exceptional ability to produce high-quality content across various domains, generative semantic communication (GSC) presents a transformative solution. In GSC systems, task-related semantic information is distilled in the form of a prompt, which is transmitted to the receiver to guide the GAI models in inference [2]. This approach reduces the required data payload and enhances the overall efficiency of information exchange. Recent research validates the effectiveness of using image captions [3], segmentation maps [4], and skeleton maps [5] as prompts to reduce the transmission payload in image transmission scenarios.
However, owing to the elementary nature of the tasks addressed, the semantically homogeneous prompts employed in the referenced studies [3, 4, 5] fall short in scenarios characterized by complex demands and high-precision requirements. For example, in vehicular networks and the IoT, achieving comprehensive and precise inference at the user end necessitates the extraction and transmission of an extensive range of semantic data from the source [6, 7]. In vehicular networks, this includes data related to vehicle behavior, traffic conditions, and driver status, each representing distinct semantic categories that require precise processing to ensure both safety and efficiency. Similarly, in IoT environments, devices must interpret a diverse set of data inputs—from sensor outputs on machine performance to user interactions and environmental changes—enabling smarter home and industrial systems to dynamically adapt to variations in usage patterns, environmental conditions, and equipment status. The necessity to handle such a broad spectrum of data poses substantial challenges in resource-constrained environments. In these settings, distilling multi-type prompts requires sequential execution of various extraction models, which can significantly prolong the overall semantic extraction process. Meanwhile, in diffusion model-based GSC systems, achieving a high inference score necessitates iterative denoising procedures [8], which consequently induce considerable inference delays. The cumulative computation delays can potentially compromise the overall task completion speed and adversely affect the user experience.
In response to prevailing challenges, this work introduces a novel communication mechanism specifically engineered to reduce task latency. Contrary to traditional approaches that necessitate complete feature extraction prior to transmission, we sequentially and independently dispatch the extracted features to the receiver, facilitating fast semantic transmission. Building upon this fundamental mechanism, we refine both the prompt engineering approach at the transmitter and the diffusion-based inference approach at the receiver. These refinements enable us to tailor the semantic extraction sequence to precisely align with the task requirements and to enhance the overall guidance effectiveness of the sequentially transmitted semantic features on the task outputs, thereby fostering semantic adaptive transmission that is specifically attuned to the task goal. By synergistically integrating the proposed mechanism with these two approaches, we advance the development of a new GSC framework, termed FAST-GSC. This framework integrates fast and adaptive semantic transmission to concurrently reduce task latency while ensuring high task performance. The contributions of this work are outlined as follows:
  • Initially, we introduce a novel communication mechanism within GSC systems, leading to the development of a fundamental framework, termed the Parallel Semantic Extraction and Inference-based GSC (PGSC) framework. This design significantly reduces task latency, particularly in complex scenarios that necessitate the extraction and transmission of multiple types of semantic features. Preliminary implementations of this framework suggest that the sequential dependencies among various units and the reception immediacy of individual units markedly influence system performance, highlighting the key directions for potential enhancement.
@@
  • On the transmitter side, we develop a novel prompt engineering approach that incorporates a temporal dimension to enhance task performance. Specifically, we employ a reinforcement learning (RL) model to discern the intrinsic sequential dependencies among the semantic units and adapt their extraction and transmission sequences according to the task. To simplify and accelerate the learning process, we implement the invalid action masking technique, which effectively excludes actions that do not contribute to the enhancement of the performance reward.
@@
  • On the receiver side, we introduce a sequential conditional denoising approach that adapts the diffusion-based inference process to the sequential transmission mechanism. Specifically, we devise a semantic difference calculation module that accurately identifies the areas within the noise space targeted by newly arrived semantic features. This identified difference is then applied to the predicted conditional noise, enhancing the guidance effectiveness of late-arriving semantic units and alleviating the stringent immediacy requirement for the reception of semantic features.
@@
  • In our experimental evaluation, we separately assess the effectiveness of the enhancements implemented at the transmitter and receiver sides. The efficacy of these enhancements is conclusively demonstrated through both quantitative analyses and visual evidence. Building upon this validation, we synergistically integrate these enhancements into the PGSC framework, culminating in the establishment of the FAST-GSC framework, which achieves a performance score comparable to the conventional GSC architecture while realizing a 52% reduction in residual latency that extends beyond the fixed inference duration.
The rest of this paper is organized as follows. In Section II, we review the related works. In Section III, we introduce the proposed PGSC framework and compare it with the conventional GSC design in respective of system latency. The motivations for refinement and the specific methods used to evolve the basic PGSC framework into the FAST-GSC framework are detailed in Section IV. The experimental evaluations are conducted in Section V, followed by the conclusions of the study and future works in Section VI.
II Related Work
II-A Conditional Diffusion Models
@@
Conditional diffusion models have emerged as a powerful class of generative models capable of producing high-quality samples across a range of data modalities. These models iteratively refine a distribution of noise to generate samples, conditioned on external information to guide the generation process towards desired characteristics. To exert a high degree of control over the generation process, a diverse array of conditions can be employed. Methods that leverage text for guidance have demonstrated remarkable capabilities across various media types [9, 10]. For delineating the contours of image objects, several studies have employed segmentation masks as conditioning inputs [11]. Additionally, the potential for manipulation through other modalities such as keypoints and depth maps has been explored, further expanding the versatility of control [12, 13].
Nonetheless, most current research introduces conditions at the start of the denoising process. Recently, [14, 15, 16] investigate the feasibility of collaborative distributed diffusion, aiming to reduce the total inference latency in multi-user scenarios by partitioning each user’s inference process into a co-denoising part executed at the edge device and a separate denoising part executed locally. However, because user-specific conditions are introduced midway through the denoising process, they may fail to effectively influence the inference results due to the already reduced noise levels. Therefore, each user’s inference performance may decrease compared to performing the entire denoising process independently. The issue of how to enhance the impact of temporally late-introduced conditions on inference results to improve overall inference performance remains an open question.
II-B Prompt Engineering
@@
Recent research in this field indicates that advanced models like DALL-E [17], Stable Diffusion [18], and GPT-4 [19] respond variably to subtle nuances in language, where even minor modifications in the prompt can lead to significantly different outputs. In light of this phenomenon, prompt engineering emerges as a crucial technique for boosting the performance of GAI models. Within this field, several methods are prominently utilized. For example, keyword-based prompting is a prompt engineering approach that incorporates specific keywords or phrases to guide the model’s attention, ensuring alignment with the desired topic or context [20]. Context-rich prompting provides detailed scenarios or background information to frame the response accurately, making the generated output more contextually appropriate and relevant [21, 22]. Iterative refinement adjusts prompts based on feedback and results, allowing for continuous improvement in the model’s performance [23].
@@
II-C Model Acceleration
As deep learning models become more complex, their high computational demands can limit practical applications, particularly on resource-constrained devices. To address this, integrating latency considerations directly into the architecture search process has emerged as an effective solution. This approach facilitates the development of models that are intrinsically optimized for enhanced performance on specific hardware configurations [24]. Additionally, model compression techniques such as pruning and quantization are increasingly recognized for their ability to reduce model size and computational complexity, thereby decreasing latency [25]. Knowledge distillation represents another promising strategy, where a compact, efficient model is trained to mimic the behavior of a larger, more intricate model [26].
Instead of solely focusing on individual models, we aim to propose to reduce the overall task latency from a holistic system perspective. Additionally, integrating the methodology developed in this study with the aforementioned model optimization techniques could further diminish system delays.
III PGSC: FRAMEWORK AND KEY COMPONENTS
In this work, we concentrate on image transmission scenarios due to their widespread prevalence and demonstrative nature. However, it is essential to note that our approach is also applicable to a variety of transmission scenarios and different data modalities. Herein, the receiver is an image renderer that produces images of varying levels of detail based on different task requirements. The transmitter, on the other hand, is an end device with limited computational capabilities and resources, responsible for capturing the key features of the source images according to the receiver’s request.
The PGSC framework can be divided into distinct modular functional components, including semantic extraction, wireless transmission, and diffusion-based inference. In the following, we first introduce semantic extraction and diffusion-based inference processes. Next, we compare the proposed PGSC framework with conventional GSC frameworks in terms of system latency and communication mechanisms.
III-A Semantic Extraction
Similar to GSC, the extraction process is dedicated to deriving semantic representations from the original source data. While existing extraction models, such as specific image captioning algorithms, are capable of distilling summary-level information from images through a single operation, they frequently generate prompts that are overly generalized and deficient in detailed specificity [27, 28]. Consequently, they may not be suitable in scenarios with high precision demands. Therefore, we deploy a series of extraction models at the transmitter, each specifically designed to accurately extract a certain type of semantic description by leveraging its specialized knowledge. The activation of these models depends on the task requests received from the receiver. Herein, we denote each extraction result as a semantic unit, and refer to the collection of all semantic units from an image as a prompt.

### v2

# FAST: Flexible and Adaptive Semantic Transmission for Resource-constrained Multi-user Generative Semantic Communication

@@ Abstract
The rapid advancement of generative artificial intelligence has spurred innovative approaches to semantic communication, giving rise to a new paradigm known as generative semantic communication (GSC). The integration of flexible cross-modal semantic extraction with generative capability–driven semantic inference substantially enhances semantic compression efficiency, demonstrating significant promise under communication resource constraints. Nonetheless, the stringent dependence on high computational power and the resulting latency continue to present major challenges, thereby limiting the feasibility of large-scale deployment. To address these challenges, we propose a novel GSC framework named FAST, which stands for flexible and adaptive semantic transmission. To accommodate limited computational resources, we propose a sequential semantic extraction method, where a temporal prompt engineering module orchestrates the distillation and transmission of key semantic units. Correspondingly, we introduce a sequential conditional denoising module at the receiver, which adapts the diffusion-based reconstruction to the progressively received input. To enhance overall task performance in multi-user semantic transmission, we propose a semantic-aware resource allocation method that optimizes bandwidth dynamically based on a joint consideration of semantic dependencies, user-level task priorities, and instantaneous channel conditions. Extensive experiments demonstrate that the proposed architecture achieves system precision comparable to conventional GSC systems while significantly reducing transmission latency and improving overall efficiency. These results confirm its enhanced potential for deployment in multi-user GSC scenarios with stringent communication and computational constraints.
@@ I Introduction
The imminent advent of sixth-generation (6G) networks is anticipated to experience an unparalleled increase in data traffic, fueled by the proliferation of cutting-edge applications such as ultra-high-definition video streaming, vehicular networks, and extensive Internet of Things (IoT) deployments. Concurrently, traditional communication systems are approaching their limit, and the scarcity of available spectrum resources is becoming more pronounced. These factors collectively present substantial challenges in ensuring the quality of user experiences for these advanced services [1]. Recognizing the capacity of generative artificial intelligence (GAI) for modal transformation and its exceptional ability to produce high-quality content across various domains, generative semantic communication (GSC) presents a transformative solution. In GSC systems, task-related semantic information is distilled in the form of a prompt, which is transmitted to the receiver to guide the GAI models in reconstruction [2]. This approach reduces the required data payload and enhances the overall efficiency of information exchange. Recent research validates the effectiveness of using image captions [3], segmentation maps [4], and skeleton maps [5] as prompts to reduce the transmission payload in image transmission scenarios.
However, conventional GSC systems [3, 4, 5] lack the adaptability for complex tasks like digital twins and virtual reality games, which are characterized by heterogeneous data sources, diverse user task criticality, and time-varying channel conditions from mobility and obstruction. For instance, a digital twin system integrates data from multiple physical sensors, with each source representing distinct semantic dimensions—from entity states and environmental interactions to system dynamics. Concurrently, system users (e.g., automated guided vehicles or inspection workers) are often mobile while also exhibiting diverse task-specific demands for critical functions such as synchronization, prediction, and optimization, each imposing distinct requirements on latency, reliability, and data fidelity. The interplay of these challenges is further compounded by the inherent computational and communication resource constraints at the network edge, creating a critical bottleneck for conventional transmission designs.
To overcome these limitations, we introduce a flexible and adaptive semantic transmission (FAST) framework for computation and communication resource-constrained multi-user GSC scenarios that efficiently handles heterogeneous data and adaptively prioritizes tasks based on their criticality. The contributions of this work are outlined as follows:
@@
  • To accommodate limited computational resources, we propose a sequential semantic extraction method. In this approach, a temporal prompt engineering module governs the execution order of extraction models and distills key semantic units for transmission in a sequential manner. Specifically, this module employs a reinforcement learning (RL) agent to discern the intrinsic sequential dependencies among the semantic units and adapt their extraction sequences according to the task. To enhance learning efficiency, we incorporate an invalid action masking technique, which prunes the action space by excluding choices that do not positively contribute to the performance reward.
@@
  • At the user end, we introduce a sequential conditional denoising module that adapts the diffusion-based reconstruction process to the sequential transmission mechanism. Specifically, we devise a semantic difference calculation subunit that accurately identifies the areas within the noise space targeted by newly arrived semantic units. This identified difference is then applied to the predicted conditional noise, enhancing the guidance effectiveness of late-arriving semantic units and alleviating the stringent immediacy requirement.
  • To enhance overall task performance in multi-user semantic transmission, we develop a semantic-aware resource allocation module. This module performs dynamic bandwidth allocation by synthesizing three critical factors: semantic dependencies among data units, user-level task priorities, and instantaneous channel conditions.
@@
  • In our experimental evaluation, we separately assess the effectiveness of the aforementioned three components. The efficacy of these enhancements is conclusively demonstrated through both quantitative analyses and visual evidence. Building upon this validation, we synergistically integrate these modules, culminating in the establishment of the FAST framework. Extensive experiments demonstrate that the proposed architecture achieves task precision comparable to conventional GSC systems, while significantly reducing transmission latency and improving overall system efficiency.
The rest of this paper is organized as follows. The background and related work on GSC are provided in Section II. The overall architecture of the proposed FAST framework is presented in Section III, and its three core components are described in Section IV. The experimental evaluations are conducted in Section V, followed by the conclusions of the study in Section VI.
II Background and Related Work
In this section, we review the background and related work on GSC. We first introduce the foundational components of GSC, which encompass semantic communication principles, conditional diffusion models, and prompt engineering techniques. Subsequently, we critically review relevant studies in GSC, and compare them with our proposed FAST framework.
II-A Fundamentals of GSC
1) Semantic Communications: Semantic communication has emerged as a paradigm shift that moves beyond the conventional goal of transmitting bits accurately toward the precise delivery of meaning [6, 7, 8, 9]. Unlike traditional systems that focus primarily on signal-level fidelity, semantic communication seeks to ensure the successful execution of a downstream task, such as classification [10, 11], reconstruction [12, 13, 14], or decision-making [15]. By discarding semantically redundant information, this approach demonstrates significant potential for extreme data compression and bandwidth savings, particularly in resource-constrained scenarios.
Existing semantic communication frameworks typically leverage a joint source-channel coding (JSCC) architecture, implemented via end-to-end trained deep neural networks, to extract and transmit task-relevant semantic features. Xie et al. [16] proposed a Transformer-based semantic communication system for text transmission. In the proposed framework, a JSCC framework was designed to extract the semantic information from texts and cope with channel noise and semantic distortion. Yang et al. [17] redesigned the vision Transformer as a backbone for image transmission. This JSCC framework was optimized to extract long-range information from images and consider the effect of the wireless channel.
Traditional SemCom faces two principal challenges: its reliance on end-to-end models results in poor interpretability, making debugging costly; furthermore, the framework exhibits limited extensibility, requiring major redesigns to handle multi-modal tasks and thus hindering its broad applicability.
@@ II-A Fundamentals of GSC
2) Conditional Diffusion Models: Conditional diffusion models have emerged as a powerful class of generative models capable of producing high-quality samples across a range of data modalities [18, 19, 20, 21]. These models iteratively refine a distribution of noise to generate samples, conditioned on external information to guide the generation process towards desired characteristics. To exert a high degree of control over the generation process, a diverse array of conditions can be employed. Methods that leverage text for guidance have demonstrated remarkable capabilities across various media types [22, 23]. For delineating the contours of image objects, several studies have employed segmentation masks as conditioning inputs [24]. Additionally, the potential for manipulation through other modalities such as keypoints and depth maps has been explored, further expanding the versatility of control [25, 26].
To address the computational challenges of deploying diffusion models on resource-constrained devices, Du et al. [27] proposed a distributed denoising framework. In this approach, the initial denoising steps are offloaded to an edge server and shared across multiple users. The resulting intermediate latent representations are then transmitted to individual user devices, which complete the remaining task-specific denoising steps. This strategy not only improves computational efficiency but also enhances privacy by allowing users to retain control over their final output. However, a key limitation is that user-specific prompts are introduced only after the shared denoising process. This delayed induction weakens their influence on the generated content, potentially leading to suboptimal reconstruction quality.
To address this limitation, Du et al. [28] enforced a constraint whereby the shared initial denoising steps were executed exclusively among users possessing semantically similar prompts. In a concurrent approach, Xie et al. [29] proposed grouping users based on task similarity and reducing divergence among the late-induced user-specific prompts, thereby enhancing the reconstruction performance at a group level. However, this method does not fundamentally address the root cause of performance degradation—the inherently limited influence of any prompt introduced after critical early denoising steps. This core limitation persists and necessitates more effective solutions.
@@
3) Prompt Engineering: Recent research in this field indicates that advanced models like DALL-E [30], Stable Diffusion [31], and GPT-4 [32] respond variably to subtle nuances in language, where even minor modifications in the prompt can lead to significantly different outputs. In light of this phenomenon, prompt engineering emerges as a crucial technique for boosting the performance of GAI models. Within this field, several methods are prominently utilized. For example, Brown et. al [33] proposed a keyword-based prompting approach that incorporates specific keywords or phrases to guide the model’s attention, ensuring alignment with the desired topic or context. Gao et. al [34] and Liu et. al [35] developed context-rich prompting which provides detailed scenarios or background information to frame the response accurately, making the generated output more contextually appropriate and relevant. Feng et. al [36] proposed iterative refinement which adjusts prompts based on feedback and results, allowing for continuous improvement in the model’s performance.
@@
II-B Review of the Existing Studies in GSC
To enhance the performance of GSC, a series of transmission designs have been developed. These approaches can be broadly categorized into those tailored for single-user and multi-user communication scenarios.
1) Single-user GSC Scenario: Within single-user scenarios, research has shown that GAI holds significant promise for enhancing key system performance metrics of semantic communications.
Grassucci et al. [37] proposed a diffusion-guided semantic communication framework that capitalizes on the powerful generative capability of diffusion models in synthesizing multimedia content. This approach achieved a significant reduction in bandwidth usage while effectively preserving semantic integrity. Yang et al. [38] developed a semantic change-driven generative semantic communication framework for remote monitoring applications. Their design incorporated a modular semantic encoder equipped with a value-of-information-based sampling function, along with a conditional denoising diffusion probabilistic model acting as the semantic decoder to accurately reconstruct remote scenes. Experimental results validated that the proposed encoder-decoder architecture exhibited significant potential for minimizing energy consumption. In their subsequent work [1], they designed an agent-assisted semantic encoder with cross-modality capabilities, which further reduced the transmission workload and enhances energy efficiency. Guo et. al [39] designed an advanced variational auto-encoder-based compression method for bandwidth-constrained GSC systems. To achieve low-latency and high-quality video transmission over bandwidth-constrained noisy wireless channels, Li et al. [40] proposed a diffusion-based goal-oriented semantic communication framework. The framework incorporated a paired semantic encoder-decoder to efficiently extract and reconstruct video information, thereby reducing transmitted data size, along with a diffusion model-based denoiser to mitigate channel-induced noise.
However, focusing solely on single-user optimization leads to a sub-optimal system-level performance, as it ignores the potential for coordinated transmission and scheduling management. Therefore, advancing multi-user GSC is imperative to unlock the full potential of semantic-aware networks and transition from isolated links to an intelligent system.
2) Multi-user GSC Scenario: The extension of GSC from single-user to multi-user architectures is not merely a scalability enhancement but a critical step toward practical deployment. The multi-user GSC scenarios involve multiple agents or devices with distinct tasks, data semantics, and quality-of-service requirements, all competing for shared and often severely limited communication and computational resources. Therefore, a well-designed transmission strategy is essential to further enhance the flexibility and adaptability of the network.
Grassucci et al. [41] proposed a channel allocation strategy based on the generative capability of diffusion models and utilized these models to reconstruct missing information at the receivers. This strategy enabled more efficient bandwidth allocation across multiple users. To address diverse yet overlapping user demands, Lu et. al [42] proposed a multi-user GSC-enhanced intent-aware semantic-splitting multiple access framework. Within this framework, the common semantic information was designated as a one-hot semantic map that was broadcast to all users, while the private semantic information was delivered as personalized text for each user. Wang et. al [43] investigated multi-user GSC system for video transmission, which comprised a base station and paired users. The base station generated and combined semantic information of several frames simultaneously requested by paired users into a signal. Users recovered their frames based on a GAI-based video frame interpolation model. To optimize transmission rates and temporal gaps between simultaneously transmitted frames, an optimization problem was formulated to maximize the system sum rate while minimizing temporal gaps. To overcome substantial bandwidth consumption and potential transmission failures of mobile AI-generated content service, Liu et. al [44] applied cross-modal GSC framework and formulated a joint semantic encoding and prompt engineering problem to optimize the bandwidth allocation among users. Ren et. al [45] developed a multi-user GSC framework leveraging pretrained Multi-modal/Vision Language Models and formulated a joint optimization problem of prompt generation offloading, communication and computation resource allocation to minimize the latency and maximize the resulting semantic quality.
The transition to multi-user semantic communication poses a significant challenge, primarily due to user task heterogeneity and stringent resource constraints—a critical gap that remains largely unaddressed in the existing literature. To situate our contribution within this context, we systematically outline the aforementioned frameworks and present a comparative analysis with our proposed FAST method in Table I.
TABLE I: A Summary and Comparison of Existing GSC Studies and Our Work
III System Model
We consider a multi-user system where computational resources for semantic extraction and communication resources for wireless transmission are shared among all users. Image transmission is adopted as a representative use case, with semantic units defined as scene entities, their textual descriptions, and inter-entity relations. This formulation enables tractable experimentation while maintaining scalability to more complex scenarios. The workflow and key components of the proposed FAST framework are illustrated in Fig. 1.
The proposed FAST framework operates through four sequential stages: 1) multi-source data collection, 2) sequential semantic extraction, 3) multi-user resource allocation, and 4) diffusion-based data reconstruction.
Initially, the edge server collects raw data from nearby sensors. Subsequently, semantic extraction models are executed in an order determined by the temporal prompt engineering module. This sequential approach is designed to efficiently utilize constrained computational resources at the edge. During the multi-user resource allocation stage, the allocated communication resources are dynamically optimized for the transmitted semantic units by the semantic-aware resource allocation module, factoring in semantic dependencies, the task importance of each user, and real-time channel conditions.
At the user end, a diffusion model initiates the reconstruction process immediately upon the arrival of the first semantic units, thereby reducing latency. This reconstruction is iteratively refined at each time step, guided by all semantic units received up to that point. As the units arrive sequentially, the guidance signal for the diffusion model evolves across its denoising steps. To accommodate this dynamic guidance, we introduce a sequential conditional denoising module, which enables the model to efficiently integrate the progressively arriving semantic information.
Figure 1: An overview of the proposed FAST-GSC framework.
IV Key Components of FAST
The temporal prompt engineering and sequential conditional denoising modules form a complementary pair: the former regulates the sequential transmission order at the transmitter, and the latter accommodates this sequence at the receiver. Designed independently yet coordinated for seamless operation, this per-user unit is elaborated first. Following this, we scale the framework to multi-user transmission, detailing the integrated semantic-aware resource allocation module for resource allocation.
IV-A Temporal Prompt Engineering
The semantic extraction process is dedicated to deriving concise semantic representations from the original source data. To accommodate heterogeneous source data distributions and diverse user demands, multiple specialized semantic extraction models can be employed to distill multi-dimensional semantic units. Herein, each extraction result is denoted as a semantic unit. For a given user, the collection of all semantic units forms an ordered prompt sequence, which we inject progressively into the diffusion process over successive denoising steps.